# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

### v2

@@ Abstract.
In Time Series Forecasting (TSF), the lookback window (the length of historical data used for prediction) is a critical hyperparameter that is often set arbitrarily, undermining the validity of model evaluations. We argue that the lookback window must be tuned on a per-task basis to ensure fair comparisons. Our empirical results show that failing to do so can invert performance rankings, particularly when comparing univariate and multivariate methods. Experiments on standard benchmarks reposition Channel-Independent (CI) models, such as PatchTST, as state-of-the-art methods. However, we reveal this superior performance is largely an artifact of weak inter-channel correlations and simplicity of patterns within these specific datasets. Using Granger causality analysis and ODE datasets (with implicit channel correlations), we demonstrate that the true strength of multivariate Channel-Dependent (CD) models emerges on datasets with strong, inherent cross-channel dependencies, where they significantly outperform CI models. We conclude with three key recommendations for improving TSF research: (i) treat the lookback window as a critical hyperparameter to be tuned, (ii) use statistical analysis of datasets to inform the choice between CI and CD architectures, and (iii) favor CD models in applications with limited data.
conference: ; ; copyright: none
1. Introduction
@@
Figure 1. Test-error (MSE) on y-axis for various lookback windows on the x-axis for the weather dataset and a forecast horizon of 96. CI denotes the channel-independent variant and CD denotes the channel-dependent variant of TSF architectures. The black dotted line shows lookback window of 96, where most methods compare performance.
@@
Time Series Forecasting (TSF) is a fundamental challenge in machine learning that drives key innovations across domains such as finance, energy, healthcare, and climate science (Godahewa et al., 2021). The core task of TSF is to predict future observations by leveraging patterns within historical time-dependent data. Nevertheless, the literature suffers from two significant experimental shortcomings that compromise the fairness and reliability of model comparisons.
First, in the multivariate TSF learning setting, the historical behavior is modeled using a fixed-length sequence of past observations—known as the lookback window—to forecast a set of future values over a given prediction horizon. Despite its fundamental role, the choice of lookback window remains surprisingly underexplored in TSF literature. Most existing Long-term Time Series Forecasting (LTSF) studies adopt a fixed lookback window (mostly 96) for all models and datasets, often without justification or transparency (Liu et al., 2024; Zeng et al., 2023; Zhang and Yan, 2023; Yi et al., 2023; Chen et al., 2023, 2025; Lin et al., 2024, 2025; Huang et al., 2025).
While this practice simplifies experimental comparisons, it masks a critical reality: the optimal lookback window can vary significantly across models and datasets, and failing to tune it can lead to misleading conclusions. As demonstrated in Figure˜1, for the weather dataset, the optimal lookback window length differs across various state-of-the-art TSF models. Consequently, performance comparisons at a single, fixed lookback window may misrepresent which models are truly state-of-the-art under realistic, well-optimized conditions. Comparable studies have been conducted by Zeng et al. (2023) on the Electricity dataset and by Tong and Yuan (2025) on the Weather dataset. However, unlike previous studies, we explicitly tune all the hyperparameters for each lookback window to generate the figure ensuring reliable comparisons.
The optimal lookback window is highly dependent on both the intrinsic properties of the dataset and the specific architecture of the forecasting model. For example, when comparing within the same weather dataset (Figure˜1), Channel Independent (CI) or univariate models tend to favor longer lookback windows (e.g., 336 or 512), enabling deeper temporal modeling per variable (Nie et al., 2023; Zeng et al., 2023; Chen et al., 2023). Whereas the Channel Dependent (CD) or multivariate models jointly model all channels and often operate with shorter lookback windows (e.g., 96), leveraging channel dependence to compensate for the reduced temporal depth (Wang et al., 2024b; Liu et al., 2024). Even within CD class of methods (i.e, across different multivariate models) certain architectures benefit more from long histories than others. For example, Crossformer-CD significantly outperforms TimeMixer-CD at a lookback window of 96, yet its performance drops below TimeMixer-CD when the lookback window is extended to 336. These findings motivate our central claim:
Tuning the lookback window as a model-specific hyperparameter is essential for fair evaluation and reliable assessment in time series forecasting.
@@
Beyond the issue of lookback tuning, a second critical limitation undermines the fairness and practical relevance of current TSF benchmarks. Most widely-used multivariate forecasting datasets such as ETTh, ETTm, Electricity, Weather, and Traffic exhibit minimal channel dependence among their variables. This lack of inter-variable relational complexity renders these benchmarks overly simplistic; even simple linear models (Zeng et al., 2023) can achieve performance comparable to sophisticated nonlinear architectures. This raises a key question:
Does channel dependence truly matter in multivariate LTSF?
@@
To answer this, we conduct an in-depth empirical analysis, combining traditional performance metrics with statistical tools such as Granger causality (Granger, 1969). Our analysis reveals that the benefits of multivariate modeling depend strongly on two factors: (1) the amount of historical data available, and (2) the degree of causal influence among channels. When inter-channel interactions are weak or datasets are large enough for univariate models to exploit long-term patterns, the added complexity of modeling channel dependence may offer limited returns. Additionally, through experiments on complex channel-dependent chaotic Ordinary Differential Equation (ODE) datasets (Gilpin, 2021), we reveal that the standard benchmarks lack realistic cross-channel interactions essential for multivariate forecasting. This simplicity disproportionately favors CI models and reduce the ability to fairly assess models designed to exploit channel dependencies as also shown in (Nie et al., 2023; Chen et al., 2023).
Based on these experimental and statistical findings, we offer the following practical recommendations for evaluating and designing multivariate time series forecasting models:
  • Channel dependence is unnecessary for top performance on standard benchmarks; tuning the lookback window enables CI models like PatchTST to match or surpass CD ones (Section˜6.1).
  • On complex datasets with strong channel correlations (supported by granger causality Section˜6.2), CD models outperform CI models (Section˜6.3).
  • Longer lookback windows generally favor univariate models, highlighting the need for careful lookback tuning to ensure fair comparisons (Section˜6.4).
2. Related Works
LTSF models typically fall into two categories: Channel-Dependent (CD), which model cross-channel interactions, and Channel-Independent (CI), which treat each channel in isolation and focus purely on temporal dynamics. CI models such as DLinear (Zeng et al., 2023) and PatchTST (Nie et al., 2023) process each feature independently and consistently achieve strong performance on widely-used benchmarks, casting doubt on the necessity of cross-channel modeling. In contrast, MLP-based CD methods like TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024a), and transformer-based architectures including iTransformer (Liu et al., 2024) and Crossformer (Zhang and Yan, 2023) aim to capture inter-channel dependencies explicitly. However, the predominantly univariate nature of standard benchmark datasets limits the ability of CD methods to demonstrate their advantages. To mitigate this issue, we incorporate evaluations on ODE-based datasets (Gilpin, 2021) designed to embed implicit channel correlations and enabling more representative comparisons on channel-dependent datasets.
The effectiveness of CI methods also appears to be amplified by longer lookback windows. For instance, DLinear adopts a lookback of 336 steps, whereas transformer-based baselines were defaulted to a lookback of 96. PatchTST, a successor to DLinear, employs a fixed lookback window of 512 or 336 and addresses the limitations of its predecessor by systematically tuning baseline methods over a broader range of lookback lengths. Nonetheless, the widespread adoption of a fixed 96-step lookback, popularized by iTransformer has constrained CI models and impeded equitable benchmarking. Subsequent studies (Tong and Yuan, 2025; Chen et al., 2025; Lin et al., 2024, 2025; Huang et al., 2025) frequently report improved performance of proposed methods by retaining short fixed windows for strong CI baselines like PatchTST, thereby hindering fair comparisons due to inconsistent tuning.
Several recent studies (Wang et al., 2024a; Nie et al., 2023; Zeng et al., 2023; Tong and Yuan, 2025) acknowledge the importance of lookback window selection in forecasting performance. However, lookback tuning is frequently confined to ablation studies, with many works failing to apply consistent tuning across both baselines and proposed methods in their main experiments. For instance, TimeMixer reports results using a unified set of hyperparameters (including a fixed lookback window) in the main results, while providing tuned results only in the appendix. Another work proposed IRPA method (Tong and Yuan, 2025) that explores various lookback lengths (480, 1920, 3600) for their method, yet retains a fixed lookback of 96 for all baselines, diminishing their comparative performance and inflating IRPA’s gains. In contrast, our evaluations perform systematic lookback window tuning for all models, uncovering genuinely optimal configurations.
3. Problem Statement
Multivariate time series is a sequence $({\bm{x}}^{l})_{l=1}^{T}\in(\mathbb{R}^{C})^{T}$, where each ${\bm{x}}^{l}\in\mathbb{R}^{C}$ represents observations from $C$ channels at time step $l$. This sequence can be represented as a matrix ${\bm{X}}^{\text{full}}\in\mathbb{R}^{C\times T}$. If $C=1$, this represents the univariate forecasting setting.
@@ 3. Problem Statement
To train forecasting models effectively, especially when $T$ is large, we partition the long sequence into multiple shorter subsequences using a sliding window approach. Each subsequence is split into two parts: an input segment (lookback window) of length $L$ and a target segment (forecast horizon) of length $H$. Specifically, a time-series forecasting dataset is composed of input-output pairs $({\bm{X}},{\bm{Y}})\in\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, where:
${\bm{X}}=({\bm{x}}^{t},{\bm{x}}^{t+1},\dots,{\bm{x}}^{t+L-1})$ is referred to as the forecasting query, and
${\bm{Y}}=({\bm{x}}^{t+L},{\bm{x}}^{t+L+1},\dots,{\bm{x}}^{t+L+H-1})$ is the corresponding forecasting answer.
@@
The goal of the time-series forecasting problem is to train a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times H}$ on a dataset $\mathcal{D}_{\operatorname{train}}\subseteq\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, constructed from many such overlapping or non-overlapping windows, sampled from the full sequence. This dataset is assumed to be drawn i.i.d. from some underlying distribution $p$. The objective is to minimize the expected loss:
(1) $\mathbb{E}_{({\bm{X}},{\bm{Y}})\sim p}\,\ell({\bm{Y}},\hat{{\bm{Y}}}({\bm{X}}))$
where $\ell:\mathbb{R}^{C\times H}\times\mathbb{R}^{C\times H}\to\mathbb{R}$ is a suitable loss function, commonly the mean squared error (MSE).
4. Background
Table 1. Average MSE over the 4 forecasting horizons for each variant of the baseline evaluated for the standard datasets. The best variant for each model is in bold and the best overall model is in blue. Note here due to running out of memory (OOM) for some experiments related to electricity and traffic, we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).

### v3

@@ Abstract
In Long-term Time Series Forecasting (LTSF), the lookback window is a critical hyperparameter often set arbitrarily, undermining the validity of model evaluations. We argue that the lookback window must be tuned on a per-task basis to ensure fair comparisons. Our empirical results show that failing to do so can invert performance rankings, particularly when comparing univariate and multivariate methods. Experiments on standard benchmarks reposition Channel-Independent (CI) models, such as PatchTST, as state-of-the-art methods. However, we reveal this superior performance is largely an artifact of weak inter-channel correlations and simplicity of patterns within these specific datasets. Using Granger causality analysis and ODE datasets (with implicit channel correlations), we demonstrate that the true strength of multivariate Channel-Dependent (CD) models emerges on datasets with strong, inherent cross-channel dependencies, where they significantly outperform CI models. We conclude with four key recommendations for improving TSF research: (i) consider the lookback window as a key hyperparameter to tune, (ii) for standard datasets, examining CI architectures is advantageous, (iii) leverage statistical analysis of datasets to guide the choice between CI and CD architectures, and (iv) prefer CD models in scenarios with limited data.
1 Introduction
@@
Figure 1: Test-error (MSE) on the y-axis for various lookback windows on the x-axis for the weather dataset and a forecast horizon of 96. CI denotes the channel-independent variant and CD denotes the channel-dependent variant. The black dotted line shows a lookback window of 96.
@@
Long-term Time Series Forecasting (LTSF) is a core machine learning task that aims to predict future observations by using patterns in historical time-dependent data in various domains such as finance, energy, healthcare, and climate science [6]. Nevertheless, the literature suffers from two significant experimental shortcomings that compromise the fairness and reliability of model comparisons.
First, LTSF models utilize a fixed-length sequence of past observations, the lookback window ($L$), for prediction. Despite its fundamental role, The lookback window is often set arbitrarily (typically $L=96$) across benchmarks without justification [15, 25, 26, 24, 3, 2, 14, 13, 9].
While simplifying comparisons, using a fixed lookback window masks a critical reality: the optimal lookback window varies across models and datasets (Figure˜1). Fixed-window comparisons may thus obscure which models truly achieve state-of-the-art performance under optimized settings. Unlike [25, 21], our study explicitly tunes all hyperparameters for every lookback window to ensure reliable evaluation.
The optimal lookback window depends on both dataset properties and model architecture. As shown in Figure˜1, Channel-Independent (CI) models typically prefer longer windows (e.g., 336–512) to capture temporal patterns [16, 25, 3], whereas Channel-Dependent (CD) models often thrive with shorter windows (e.g., 96) by leveraging cross-channel dependencies [23, 15]. Even within the CD class, benefits vary: Crossformer-CD outperforms TimeMixer-CD at $L=96$, but falls behind at $L=336$. These findings motivate our first claim: Tuning the lookback window as a model-specific hyperparameter is essential for fair evaluation and reliable assessment in time series forecasting.
@@
The second limitation concerns the practical relevance of current LTSF benchmarks. Standard multivariate datasets—such as ETT, Electricity, Weather, and Traffic—exhibit very limited dependence among channels. This lack of inter-variable complexity renders these benchmarks overly simplistic, raising a critical question: Does channel dependence truly matter in multivariate LTSF?
Table 1: Frequency of CI and CD wins across models and datasets. ’ODE-6’ denotes the 6 ODE datasets with intrinsic channel correlation; ’Std.-7’ represents the 7 standard datasets.
Model ODE-6 Std.-7
CI CD CI CD
PatchTST 4 2 6 1
TSMixer 0 6 7 0
Crossformer 0 6 4 3
TimeMixer 2 4 4 3
Totals 6 18 21 7
@@
To address this, we conduct an empirical analysis combining traditional performance metrics with Granger causality [7]. Our results reveal that multivariate modeling benefits depend on: (1) available historical data and (2) the degree of causal influence among channels. When inter-channel interactions are weak, modeling channel dependence offers limited returns. Consequently, we utilize complex, chaotic Ordinary Differential Equation (ODE) datasets [5], which demonstrate that standard benchmarks fail to capture the cross-channel interactions crucial for multivariate forecasting.
As shown in Table˜1, CD variants consistently outperform CI counterparts on highly coupled ODE datasets. Conversely, standard datasets’ simplicity disproportionately favors CI models, limiting the fair evaluation of methods designed for channel dependencies [16, 3].
Our main contributions are as follows:
  • Lookback Evaluation: We show that tuning the lookback window is a critical, model-specific requirement; neglecting it introduces significant baseline bias.
  • Benchmark Simplification: We reveal that standard LTSF benchmarks lack sufficient channel dependence, unfairly handicapping CD models.
  • Causal Analysis: Using Granger causality and ODE datasets, we prove CD models excel only when complex, verifiable inter-channel interactions are present.
  • Guidelines: We provide actionable recommendations for choosing between CI and CD architectures based on lookback and causal complexity.
2 Related Work
LTSF models typically fall into two categories: Channel-Dependent (CD) and Channel-Independent (CI). CI models like DLinear [25] and PatchTST [16] emphasize temporal dynamics, often outperforming CD models such as TSMixer [3], iTransformer [15], and Crossformer [26]. Recent advancements in CD modeling, including TimeFilter’s [8] graph filtration and DUET’s [18] hierarchical clustering, attempt to capture complex spatial dependencies, yet the predominantly univariate nature of common benchmarks often masks their potential. Recently, a surge in using Ordinary Differential Equations (ODEs) has generated rigorous, high-causality benchmarks, exemplified by Physiome [12] and chaotic ODE collections [5]. These provide mathematically-defined channel correlations, offering a more robust multivariate testbed than standard datasets.
A critical factor in model evaluation is the lookback window length ($L$). While CI methods typically benefit from $L\in\{336,512\}$, many studies [21, 2, 13, 9] fix $L=96$ for baselines. This practice introduces evaluation bias; for instance, IRPA [21] explores variable lookbacks for its own method but maintains $L=96$ for baselines, potentially overstating gains. Although some works acknowledge lookback sensitivity [22, 16], systematic tuning is often relegated to secondary ablations. While comprehensive benchmarks like TFB [17] advocate for fair evaluation across diverse datasets, they do not address the compounding issues of lookback bias and dataset coupling. In contrast, our study systematically tunes $L$ for all models and utilizes ODE-based evaluations to ensure architectural comparisons are based on truly optimal configurations.
3 Problem Statement
@@
To train forecasting models effectively, especially when $T$ is large, we partition the long sequence into multiple shorter subsequences using a sliding window approach. Each subsequence is split into an input segment of length $L$ (representing the lookback window) and a target segment of length $H$ (the forecast horizon). A forecasting dataset thus consists of pairs $(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, where $\mathbf{X}=(\mathbf{x}^{t},\dots,\mathbf{x}^{t+L-1})$ is the forecasting query and $\mathbf{Y}=(\mathbf{x}^{t+L},\dots,\mathbf{x}^{t+L+H-1})$ is the corresponding forecasting answer.
@@
The goal of the time-series forecasting task is to learn a model $\hat{\mathbf{Y}}:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times H}$ using a dataset $\mathcal{D}_{\text{train}}\subseteq\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, constructed from many such windows sampled from the full sequence. The dataset is assumed to be drawn i.i.d. from an underlying distribution $p$. The objective is to minimize the expected loss, $\mathbb{E}_{(\mathbf{X},\mathbf{Y})\sim p}\,\ell(\mathbf{Y},\hat{\mathbf{Y}}(\mathbf{X}))$, where $\ell$ is typically the mean squared error (MSE). While the underlying time series is intrinsically dependent, this formulation treats the windowed pairs $(\mathbf{X},\mathbf{Y})$ as realizations of a stationary process for the purpose of supervised optimization. This allows the model $\hat{\mathbf{Y}}$ to explicitly focus on capturing the temporal and inter-channel dependencies within each local windowed segment.
4 Background
Fixing Lookback Windows:
The problem formulation in Section˜3 does not strictly require the forecasting query $\mathbf{X}$ to use a fixed lookback length $L$. However, most recent LTSF models adopt a fixed window (commonly $L=96$) during training, primarily for implementation convenience or adherence to benchmarking standards. As argued previously, the lookback window should be treated as a model-specific hyperparameter, integral to the model design.
Channel Independence:
Recent studies [16, 3] demonstrate that on widely used benchmarks, CI models often suffice for accurate forecasting. This suggests that cross-channel interactions may be unnecessary for certain tasks. Formally, a model $\hat{\mathbf{Y}}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times H}$ is channel-independent if there exists a univariate predictor $f:\mathbb{R}^{L}\to\mathbb{R}^{H}$ such that for each channel $c\in\{1,\dots,C\}$:
$\hat{\mathbf{Y}}(\mathbf{X})_{c}=f(\mathbf{x}_{c,:})\quad\forall c\in\{1,\dots,C\}.$ (1)
Granger Causality.
Granger causality [7] evaluates whether a channel $c_{2}$ provides predictive information for $c_{1}$ beyond $c_{1}$’s own history. We compare the univariate sum of squared residuals ($SSR_{u}$) from a restricted model (CI assumption) and the multivariate ($SSR_{mv}$) from an unrestricted model (CD assumption) incorporating $c_{2}$. Using $x_{c,t}$ to denote the observation of channel $c$ at time $t$, the residuals are:
$SSR_{u}$ $=\sum_{t=L+1}^{L+H}\bigg(x_{c_{1},t}-a_{0}-\sum_{i=1}^{L}a_{i}x_{c_{1},t-i}\bigg)^{2},$ (2)
$SSR_{mv}$ $=\sum_{t=L+1}^{L+H}\bigg(x_{c_{1},t}-a_{0}-\sum_{i=1}^{L}a_{i}x_{c_{1},t-i}-\sum_{i=1}^{L}b_{i}x_{c_{2},t-i}\bigg)^{2}.$ (3)
where $a_{0}$ is the intercept, and $a_{i},b_{i}$ are learned coefficients for the lagged values of $c_{1}$ and $c_{2}$ respectively.
The $F$-statistic measures the significance of the error reduction:
$F=\frac{(SSR_{u}-SSR_{mv})/L}{SSR_{mv}/(N-k_{mv})},$ (4)
where $L$ is the number of added parameters (lags of $c_{2}$), $N$ the number of observations, and $k_{mv}$ the total parameters in the multivariate model. A high $F$-statistic indicates that $c_{2}$ Granger-causes $c_{1}$.
5 General Experimental Setup
5.1 Models
We evaluate several baselines: MLP-based (TSMixer [3], TimeMixer [23]), attention-based (PatchTST [16], iTransformer [15], Crossformer [26]), and DLinear [25]. We use DLinear-CI and iTransformer-CD to prevent architectural overlap with TSMixer; others are tested in both variants. Details are in our repository. [^1] https://github.com/IbramMedhat/How-Biased-is-Time-Series-Forecasting- and the appendix.
5.2 Datasets
The Standard Datasets:
The first category includes widely used benchmarks from recent state-of-the-art studies and surveys [16, 15, 23, 3, 26, 25]: Weather, Electricity, Traffic, and the Electricity Transformer Temperature (ETT) datasets. These benchmarks gained popularity following their adoption in the Informer paper [27].
The Chaotic ODE Benchmark:
As a second dataset collection, we consider the benchmark proposed by [5]. This benchmark was introduced as a time-series collection that features complex time series dynamics with clear real world data properties, providing a challenging and realistic setting for CD time-series forecasting. A demonstration for the DoublePendulum dataset and how it simulates a recurrent patterns across different channels can be seen in Figure˜2. This figure particularly emphasizes periods of synchronization in the angular velocities of the two pendulums during the final 200 timesteps, indicative of inter-channel correlation.
Figure 2: Time series visualization of the Double Pendulum ODE system. Notably, the bottom two channels (angular velocities of the two pendulums) exhibit synchronized behavior in the last 200 time steps, demonstrating time-varying inter-channel correlation.
We select a subset of datasets from this benchmark based on two distinctive criteria: (i) The datasets follow complex trajectories that challenge most time series models and (ii) they are derived directly from the underlying ODEs, ensuring a clear mathematical correlation between variables. For criterion (i), we use the Largest Lyapunov exponent [5] to quantify dataset complexity, selecting datasets with high values to ensure forecasting remains non-trivial task that cannot be solved easily based on one channel information. To illustrate criterion (ii), we show below a simplified form of the Double Pendulum system [4], assuming equal masses and rod lengths. The coupling between angular velocities $\frac{d\theta_{1}(t)}{dt},\frac{d\theta_{2}(t)}{dt}$ and angles $\theta_{1}(t),\theta_{2}(t)$ at time $t$ from Equations˜5 and 6 is evident, with parameters $g$ (gravity) and $l$ (rod length) incorporated. Such systems are essential in modeling real-world physics processes, including robotics [10].
$\frac{d^{2}\theta_{1}(t)}{dt^{2}}=-\frac{2g}{l}\theta_{1}(t)+\frac{g}{l}\theta_{2}(t)$ (5)
$\frac{d^{2}\theta_{2}(t)}{dt^{2}}=\frac{2g}{l}\theta_{1}(t)-\frac{2g}{l}\theta_{2}(t)$ (6)
5.3 Hyperparameter Tuning
We ensure fair evaluation via Bayesian optimization (Optuna [1]) over 20 trials per model–dataset–horizon. The lookback window $L$ is tuned across $\{96,\allowbreak 192,\allowbreak 336,\allowbreak 512,\allowbreak 720\}$, with the best variant selected via validation MSE. We prioritize MSE as the primary metric to remain consistent with LTSF benchmarks and to penalize large outliers common in complex series. Crucially, we address data-loading biases noted in [17] by setting drop_last=False during testing. This ensures no trailing points are discarded, providing a complete, unbiased evaluation across the entire distribution and confirming that lookback-CD interactions are not artifacts of data truncation.
5.4 Granger Causality
We perform Granger causality analysis following [7] with two preprocessing steps: (i) assessing stationarity and applying iterative differencing where necessary, and (ii) removing one channel from any pair with a Pearson correlation above 0.95 to reduce redundancy. Following the $F$-test (Section˜4), we report the average $F$-score and the percentage of pairs rejecting the null hypothesis.
6 Results
6.1 Evaluation of Standard Datasets
We reproduce existing time-series forecasting models under our evaluation setup, performing extensive hyperparameter tuning (including lookback window) and averaging results over five random seeds. Reported scores in Table˜2 are averaged across horizons 96, 192, 336, and 720.
PatchTST remains the state-of-the-art model in our fair evaluation. Its CI variant achieves the best performance on three of seven datasets and ranks second or third on another three. Only on the Weather dataset does it fall outside the top three, likely due to stronger inter-channel dependencies indicated by Granger causality. Overall, PatchTST attains an average rank of 2.14, a notable margin (0.57) ahead of the second-best model, CI TSMixer. Linear models continue to perform competitively. DLinear, for instance, achieves an average rank of 5.43 out of 10 across the six standard datasets and ranks second on ETTh1, slightly behind PatchTST.
Table 2: Mean MSE across four horizons. Bold: best per model; blue: overall best. (*): sourced from [22]. OOM: Out-of-memory.